OpenStudy (anonymous):

I need help with this, write using a+bi equation form 6/9-7i

OpenStudy (anonymous):

Do you mean 6/(9-7i)?

OpenStudy (anonymous):

multiply it by its conjugate and it should all work out

OpenStudy (anonymous):

\[\frac{6}{9-7i} =\frac{6}{9-7i} \times\frac{9+7i}{9+7i} \] \[ =\frac{6(9+7i)}{(9-7i)(9+7i)}\] \[ =\frac{54+42i}{81+49}\] \[ =\frac{54+42i}{130}\] \[ =\frac{27}{65}+\frac{21}{65}i\]

OpenStudy (anonymous):

thank u!!!!